40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Engineering group Lab elog  Not logged in ELOG logo
Message ID: 262     Entry time: Tue Jan 25 18:25:22 2022
Author: Stephen 
Type: Progress 
Category: RTS for COC 
Subject: Anneal of SN1009 at 400C for 10 hours 

SN1009 (80 mm witness sample) was annealed between 25 and 26 January 2022 - used the large furnace in Downs 221, atmosphere was air, used glass petri dishes to protect optic (from Gabriele and CRiMe). Start time was 17:48 Pacific, and sample was removed at ~17:30 Pacific on the next day.

Controller parameters:

Ramp up rate = 100 °C per hour
Hold temp = 400 °C
Hold time = 10 hours
Ramp down rate = 100 °C per hour

Additional annealing run after ENG_Labs/260 at higher temperature to hopefully yield improvements to absorption.

Witness RTD observed a 421.1 °C overshoot and 14.5 °C offset from controller RTD. These parameters were consistent with the elog entries ENG_Labs/253, ENG_Labs/254, and ENG_Labs/260. See attached image of the temperature profile (.xlsx file is source document).

Optic is now with Liyuan in RTS for recharacterization after anneal.

Attachment 1: anneal_sn1009_400C_20220126.png  40 kB  Uploaded Thu Jan 27 10:25:31 2022  | Hide | Hide all
anneal_sn1009_400C_20220126.png
Attachment 2: TMC01004_sn1009_400C_20210126.xlsx  137 kB  Uploaded Thu Jan 27 10:26:14 2022
ELOG V3.1.3-